current of 200 mA t is placed in mge field of magnitude 500 T that is dieced ersllel to e plane of the loop a gnted in the i Wha is e magnetic field 281 N neNm (e
Answer #1

5) turn ratio = N2/N1 = V2/ V1

= 1610 / 430 = V2/120

V2 = 449.3 V = 450 V

V2/V1 = I1/I2

450 / 120 = I1/12

I1 = 45 A